Evaluate the following


"n\sum i=1 (3/2^(i-1))"


1
Expert's answer
2022-04-16T05:38:05-0400

The aim is to compute the sum: "n\sum_{i=0}^ni". The sum "\sum_{i=0}^ni" is an arithmetic progression. It is a well-known formula. "n\sum_{i=0}^ni=n(0+1+2+\ldots+n)=n\left(\frac{n(n+1)}{2}\right)=\frac{n^2(n+1)}{2}."